OpenStudy (anonymous):

Someone please help me!!!!!! Solve the equation medal rewarded The square root of x = the square root of x-5, plus one

OpenStudy (anonymous):

\[\sqrt{x} = \sqrt{x-5} + 1\]

OpenStudy (anonymous):

Both sides of the equation are positive, so you can square both sides without getting into trouble. Try that.

OpenStudy (anonymous):

squaring both sides - x = x-5 + 1 + 2 sqrt(x-5) => 4 = 2 sqrt(x-5) 2 = sqrt(x-5) squaring both sides again - 4 = x-5 x = 9.,